### v2

@@ 1 Introduction
First passage time models (also known as first hitting time models) are a fundamental class of stochastic process models that describe the time it takes for a random process to exit a region of interest. These models are widely used in various fields, including physics, finance, biology, and cognitive science, to analyze phenomena such as the time until a stock hits a particular price level, the time it takes for a neuron to fire, a survival time of a transplant patient, or the time to reach a decision barrier that triggers a choice in psychological experiments [3, 36, 18, 32]. Accurately computing the density of the first passage time is crucial for the practical application of these models, such as in scientific hypothesis testing and parameter estimation. Since analytical solutions for many complex real-world first passage time models are unavailable, effective numerical methods are an indispensable part of the computational toolkit for applying these models. In this paper, we introduce a fast and flexible method for computing the first passage time density in a large class of models arising in cognitive neuroscience, mathematical psychology, and economics.
@@
where $W(t)$ is a standard Brownian motion. The drift rate $\mu$ quantifies the strength of the evidence that favors A over B and affects the joint distribution of the choice and the response time. Models such as this are often called drift diffusion models (DDMs) [45, 14] and written in differential form, such as,
@@
This basic DDM can be modified in a variety of ways to better model experimental data and test theories about the biological processes underlying decision making. For instance, the boundaries ($\pm a$) can be asymmetric and vary over time to reflect changing pressures to make a decision; the starting position $X(0)$ can be nonzero or random, representing an initial decision bias; the rates of drifting towards the boundaries can depend on the choices and vary over time; and the random process $X(t)$ can have more complicated stochastic dynamics than Brownian motion [14]. We adopt the terminology from [50] and call these models generalized drift diffusion models (GDDMs). Simple, computationally tractable versions of GDDMs have found widespread use in cognitive neuroscience and mathematical psychology [44, 45, 17, 11, 64]. While more complex versions of GDDMs have been proposed to encode a broader range of scientific hypotheses, their quantitative assessment and applications are currently limited due to computational challenges in parameter estimation [14]. In particular, there do not exist fast algorithms for calculating the likelihood function of the parameters (i.e., the probability density of the hitting time on the boundary reached) in some of the more complex GDDMs that are currently of scientific interest.
@@
Computational speed is crucial in many cases for the following reason: As noted in [32], first passage time models commonly incorporate covariates that may vary across individuals and even within a single trial. This is particularly relevant in computational cognitive neuroscience, where the precise specification of the GDDM might change every time a subject makes a choice, i.e., in each experimental trial. An experiment with, say, $50$ subjects, each completing $200$ trials, might require evaluating the likelihoods of $10^{4}$ different versions of the GDDM to get the overall dataset likelihood. For instance, in equation (1) the drift rate $\mu$ might be different on every trial if the choices presented to subjects are different on every trial. The problem becomes even more difficult when the trial-specific drift rate can vary dynamically within a trial.
@@
For a specific example, our original motivation to study this problem stems from attentional drift diffusion models (aDDMs)[29], which have been proposed as a generative model for investigating how visual attention influences decision making. In aDDMs, the drift rate $\mu(t)$ in each trial is allowed to covary on a moment-by-moment basis, depending on the subject’s eye movements toward one choice option or another, which are also recorded in the experimental setup. The entire shape of $\mu(t)$ changes dramatically on every trial. This model is a generic example of the general class of GDDMs where instantaneous within-trial dynamics can be based on exogenous covariates such as brain activity or eye gaze position, hence complicating the numerical procedures for computing the first passage time densities and further applications. On top of this, statistical inference tasks would still require repeated evaluations of the total dataset likelihood at different parameter values, both in frequentist and Bayesian frameworks. In particular, hierarchical Bayesian inference is widely employed in cognitive science to estimate both group- and subject-level parameters simultaneously [63, 14]. The heavy reliance on iterative Markov chain Monte Carlo (MCMC) methods in this setting further intensifies the need for fast likelihood evaluations.
@@
for possibly random $X(0)$ and for functions $\mu$ and $\sigma$ satisfying the so-called Cherkasov condition [7, 47]. Loosely speaking, the Cherkasov condition means that the complicated SDE in equation (2) can be transformed into a simpler SDE more akin to equation (1). A special case that includes many cases of practical interest, such as the aDDM, is when $\mu$ and $\sigma$ depend only on $t$ and not on X(t)X(t) [^1] In the cognitive science context, the terminology “DDM” is typically adopted to denote these models that directly integrate noisy accumulating evidence. The more general model in equation (2), including some examples shown later in this paper (e.g., the Ornstein-Uhlenbeck model), fall within the GDDM framework.. There exist numerical approaches and software for approximating hitting time densities in models even more general than equation (2), but in many cases they are currently too slow for practical use in real-world psychological experiments.
@@
Our approach shares some conceptual similarities to a line of previous work [60, 39, 43, 61, 26]. Both those methods and ours transform the problem into a canonical representation, subdivide time into discrete stages that allow for analytical approximations, compute the final approximation through integration over these stages, and then map the result back to the original problem. However, none of these earlier approaches are directly applicable to our setting or capable of statistical computation at the scale of the present work (see Section 3). In contrast, by leveraging accurate analytical results and numerical integration in a sequential manner, our semi-analytical-numerical approach achieves substantially higher computational efficiency. As shown in Section 5.4.1, it only requires microseconds for our method to evaluate a single likelihood value in a realistic aDDM example. These innovations and efficiency gains are crucial for practical likelihood-based inference in modern experimental settings.
@@
The remainder of this article is structured as follows. Section 2 begins with a mathematical description of the GDDM. Section 3 summarizes previous work related to this paper, including other efforts to address the computational challenges in first passage time models. In Section 4, we introduce our proposed method for computing the FPTDs of certain GDDMs, starting with analytical results for simple DDMs, then combining them sequentially based on Markovian properties. We also apply variable transformations and piecewise linear approximations to extend the method to more complex GDDMs, including those that allow for position-dependent drift and diffusion terms as well as dynamic decision boundaries [57, 14, 19]. Section 5 shows several numerical examples that demonstrate the speed and accuracy of our method for both likelihood computation and statistical inference tasks. Section 6 summarizes the present work, including its applicability and suggestions for future work. It is worth mentioning that while our examples are drawn from cognitive neuroscience and mathematical psychology, the method we propose is equally applicable to first passage time models in other fields. For instance, it could be extended to financial applications such as pricing double barrier options [33], as well as other areas where first passage times play a critical role.
@@ 2 Drift Diffusion Model
as the first passage time densities (FPTDs) on the upper boundary and lower boundary [^2] Note that “first passage time density on the upper (lower) boundary” is sometimes interpreted as the density of $\tau_{u}=\inf\{t>0:X(t)\geq u(t)\}(\tau_{\ell}=\inf\{t>0:X(t)\leq\ell(t)\})$ when only considering one upper boundary $u(t)$ (lower boundary $\ell(t)$). In this case $\tau=\tau_{u}\wedge\tau_{\ell}$. These quantities represent different aspects compared to our definition of the FPTD on the upper and lower boundaries and should not be confused., respectively, implicitly assuming throughout sufficient regularity conditions on the model for these densities to exist (see the sufficient conditions for density existence in [16, 54]). The marginal density of $\tau$ is thus given by $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathbb{P}(\tau\leq t)=f_{u}(t)+f_{\ell}(t)$. With this notation, we can write the joint probability density of $(\tau,C)$ as
@@
As mentioned in the Introduction, we are motivated by practical problems for which we need to evaluate many different joint densities (3). We define $\boldsymbol{\Xi}=\{\mu(\cdot),\sigma(\cdot),u(\cdot),\ell(\cdot),q_{0}(\cdot)\}$ to be the configuration of a GDDM, where $q_{0}$ is a representation for the distribution of $X(0)$, usually a probability density function or a point mass at a deterministic starting point $x_{0}$. In statistical applications, $\boldsymbol{\Xi}$ will depend on unknown parameters and known covariates, some of which might vary within and across each trial. If the $i$-th trial results in the observation $(\tau_{i},C_{i})$, then the contribution to the likelihood for that trial is $f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i})$. This is simply the FPTD on the boundary $C_{i}$ evaluated at time $\tau_{i}$ using the GDDM configuration $\boldsymbol{\Xi}_{i}$ for the $i$-th trial. In experiments with a time limit, some trials might result in non-response, in which case the contribution to the likelihood is the NPP for that GDDM configuration, say $Q_{i}=Q(\boldsymbol{\Xi}_{i})$. Consequently, denoting $R_{i}=\mathbbm{1}\{\text{a response on trial $i$}\}$, the likelihood function is then given by [^3] The statistical model for $\tau$ here is a mixed distribution consisting of a continuous part supported on $\tau\in(0,T]$ (response) and a point mass at $\tau>T$ (no response). Mixed likelihoods of this form are common in survival analysis, for example see the Section 2.2 of [31].
@@ 3 Related Work
Analytic or semi-analytic methods. For certain simple DDMs, analytic formulas for the first passage time densities or distributions can be derived, often in the form of infinite series. These methods are fast and accurate, but are typically applicable only for very specific cases. For example, they are available for the standard DDM where all parameters $\mu,\sigma,x_{0},u,\ell$ are constant [38, 4, 21]. More recently, [53] used martingale theory to derive closed-form formulas that extend certain theoretical analyses to GDDMs.
@@
The approach in this paper is based on numerically extending these fast special cases to much broader classes of GDDMs. Along this line, earlier works [60, 39, 43, 61] computed NPPs, while [26], [25], and the present work target FPTDs, although [26] and [25] restrict attention to the special case of a one-sided boundary. The FPTD case is more challenging but also more useful for statistical inference since it directly gives the likelihood. A key limitation of many prior approaches [60, 43, 61, 25, 26], is that they rely on Monte Carlo integration, which is infeasible for large-scale real-world applications.
@@
Numerical solutions of governing equations. These methods exploit the known connections between diffusion processes and partial differential equations (PDEs), called Kolmogorov equations (see the Chapter 8 of [62] for an introduction, and Chapter 5.7 of [27] for a mathematically rigorous treatment). Both the Kolmogorov forward equation (KFE, also known as Fokker-Planck equation in physics) [50, 10] and the Kolmogorov backward equation (KBE) [59, 58] have been exploited to derive the FPTDs. Moreover, the first passage time problem can be formulated into Volterra integral equations of renewal type [51, 42], where applications to DDMs [52, 40] have also been explored. In these works, numerical methods, particularly finite difference methods [55], have been developed to solve these equations, enabling the efficient computation of FPTDs in certain cases. Other numerical methods like discontinuous Galerkin methods have also been considered, see [41] and references therein. However, these approaches remain computationally expensive, especially in settings where the GDDM configuration differs across trials, including but not limited to cases with trial-specific covariates that vary dynamically over time. For a comprehensive review of these methods, we refer the readers to [48].
@@
Stochastic simulation. Methods that can sample from the hitting time distribution of the target GDDM can be used to approximate FPTDs via density estimation. Efficient methods for exact sampling are challenging; see [23, 24] and the references therein. If the goal is statistical inference, then these sampling algorithms can be incorporated into a variety of simulation-based inference (SBI) procedures; see [8] for review and [13] for an SBI example with DDMs. Nevertheless, the Monte Carlo methods underlying SBI are computationally intensive and can become impractical for large datasets, particularly when $\Xi$ involves high-dimensional parameters or covariates that vary dynamically over time and across trials.
@@ 4.1 Single-stage Model
[22] derives the following analytical formulas [^4] Another work along this direction is [1]. We adopt the results in [22] since they are simpler and were shown to agree with those of [1]. Equation (4.1), (4.1), and (4.1) have been slightly rewritten for use in this paper, compared to their original form in [22]. for FPTDs and NPDs in the single-stage model for the special case when $x_{0}=0$ and $\sigma=1$: We define $\overline{a}=\frac{a_{1}+a_{2}}{2},\ \overline{b}=\frac{b_{1}+b_{2}}{2},\ c=a_{1}-a_{2},\ b=\frac{b_{2}-b_{1}}{2}$, then
@@ 4.5 Transformations
A particularly simple $\widetilde{X}$ is standard Brownian motion, perhaps with a random starting point, i.e., equation (2) with $\mu(x,t)\equiv 0$ and $\sigma(x,t)\equiv 1$. A wide range of commonly used SDEs can be transformed into standard Brownian motion in this way, including geometric Brownian motion, the Ornstein-Uhlenbeck (O-U) process, and the Brownian bridge, among others. The question of whether a diffusion process can be transformed to a Brownian motion through transformation of the type in equation (13) was raised by Kolmogorov in [28]. Cherkasov gave a constructive proof of the necessary and sufficient condition for the existence of such a transformation in [7], which is now commonly referred as the Cherkasov condition in the literature. This was recast into a simpler form in [47], and was later used in [46] to obtain the one-sided first passage time density through the integral equation method. We present the Cherkasov condition in [47] in Appendix A.2, and instead of focusing on the condition itself, we provide specific examples of the transformation from diffusion processes to Brownian motion that are common in practical applications.
@@
where $\theta\neq 0,\lambda\in\mathbb{R},\sigma>0$ are constants. When $\theta>0$ it models the “mean-reverting” phenomenon as the process tends towards $\lambda$ over time, and it will converge to its stationary distribution $\mathcal{N}(\lambda,\frac{\sigma^{2}}{2\theta})$ as $t$ goes to infinity. This process is used, for instance, in physics to model the velocity of a particle under friction, in finance to capture mean-reverting behavior of interest rates and volatility [18], and in cognitive science to model evidence accumulation and neural dynamics [57, 6].
The solution of (18) is given by [18]
@@
The stochastic integral can be viewed as a time-changed Brownian motion [12]:

### v3

@@ 1 Introduction
First passage time models (also known as first hitting time models) are a fundamental class of stochastic process models that describe the time it takes for a random process to exit a region of interest. These models are widely used in various fields, including physics, finance, biology, and cognitive science, to analyze phenomena such as the time until a stock hits a particular price level, the time it takes for a neuron to fire, a survival time of a transplant patient, or the time to reach a decision barrier that triggers a choice in psychological experiments [bhattacharya2021random, metzler2014first, gardiner2009stochastic, lee2006threshold]. Accurately computing the density of the first passage time is crucial for the practical application of these models, such as in scientific hypothesis testing and parameter estimation. Since analytical solutions for many complex real-world first passage time models are unavailable, effective numerical methods are an indispensable part of the computational toolkit for applying these models. In this paper, we introduce a fast and flexible method for computing the first passage time density in a large class of models arising in cognitive neuroscience, mathematical psychology, and economics.
@@
where $W(t)$ is a standard Brownian motion. The drift rate $\mu$ quantifies the strength of the evidence that favors A over B and affects the joint distribution of the choice and the response time. Models such as this are often called drift diffusion models (DDMs) [ratcliff2008, fengler2022] and written in differential form, such as,
@@
This basic DDM can be modified in a variety of ways to better model experimental data and test theories about the biological processes underlying decision making. For instance, the boundaries ($\pm a$) can be asymmetric and vary over time to reflect changing pressures to make a decision; the starting position $X(0)$ can be nonzero or random, representing an initial decision bias; the rates of drifting towards the boundaries can depend on the choices and vary over time; and the random process $X(t)$ can have more complicated stochastic dynamics than Brownian motion [fengler2022]. We adopt the terminology from [shinn2020flexible] and call these models generalized drift diffusion models (GDDMs). Simple, computationally tractable versions of GDDMs have found widespread use in cognitive neuroscience and mathematical psychology [ratcliff1978, ratcliff2008, frank2015, doi2020, yartsev2018]. While more complex versions of GDDMs have been proposed to encode a broader range of scientific hypotheses, their quantitative assessment and applications are currently limited due to computational challenges in parameter estimation [fengler2022]. In particular, there do not exist fast algorithms for calculating the likelihood function of the parameters (i.e., the probability density of the hitting time on the boundary reached) in some of the more complex GDDMs that are currently of scientific interest.
@@
Computational speed is crucial in many cases for the following reason: As noted in [lee2006threshold], first passage time models commonly incorporate covariates that may vary across individuals and even within a single trial. This is particularly relevant in computational cognitive neuroscience, where the precise specification of the GDDM might change every time a subject makes a choice, i.e., in each experimental trial. An experiment with, say, $50$ subjects, each completing $200$ trials, might require evaluating the likelihoods of $10^{4}$ different versions of the GDDM to get the overall dataset likelihood. For instance, in equation (1) the drift rate $\mu$ might be different on every trial if the choices presented to subjects are different on every trial. The problem becomes even more difficult when the trial-specific drift rate can vary dynamically within a trial.
@@
For a specific example, our original motivation to study this problem stems from attentional drift diffusion models (aDDMs)[krajbich2010visual], which have been proposed as a generative model for investigating how visual attention influences decision making. In aDDMs, the drift rate $\mu(t)$ in each trial is allowed to covary on a moment-by-moment basis, depending on the subject’s eye movements toward one choice option or another, which are also recorded in the experimental setup. The entire shape of $\mu(t)$ changes dramatically on every trial. This model is a generic example of the general class of GDDMs where instantaneous within-trial dynamics can be based on exogenous covariates such as brain activity or eye gaze position, hence complicating the numerical procedures for computing the first passage time densities and further applications. On top of this, statistical inference tasks would still require repeated evaluations of the total dataset likelihood at different parameter values, both in frequentist and Bayesian frameworks. In particular, hierarchical Bayesian inference is widely employed in cognitive science to estimate both group- and subject-level parameters simultaneously [wiecki2013hddm, fengler2022]. The heavy reliance on iterative Markov chain Monte Carlo (MCMC) methods in this setting further intensifies the need for fast likelihood evaluations.
@@
for possibly random $X(0)$ and for functions $\mu$ and $\sigma$ satisfying the so-called Cherkasov condition [cherkasov1957transformation, ricciardi1976transformation]. Loosely speaking, the Cherkasov condition means that the complicated SDE in equation (2) can be transformed into a simpler SDE more akin to equation (1). A special case that includes many cases of practical interest, such as the aDDM, is when $\mu$ and $\sigma$ depend only on $t$ and not on X(t)X(t) [^1] In the cognitive science context, the terminology “DDM” is typically adopted to denote these models that directly integrate noisy accumulating evidence. The more general model in equation (2), including some examples shown later in this paper (e.g., the Ornstein-Uhlenbeck model), fall within the GDDM framework.. There exist numerical approaches and software for approximating hitting time densities in models even more general than equation (2), but in many cases they are currently too slow for practical use in real-world psychological experiments.
@@
Our approach shares some conceptual similarities to a line of previous work [wang1997boundary, novikov1999approximations, potzelberger2001boundary, wang2007crossing, jin2017first]. Both those methods and ours transform the problem into a canonical representation, subdivide time into discrete stages that allow for analytical approximations, compute the final approximation through integration over these stages, and then map the result back to the original problem. However, none of these earlier approaches are directly applicable to our setting or capable of statistical computation at the scale of the present work (see Section 3). In contrast, by leveraging accurate analytical results and numerical integration in a sequential manner, our semi-analytical-numerical approach achieves substantially higher computational efficiency. As shown in Section 5.4.1, it only requires microseconds for our method to evaluate a single likelihood value in a realistic aDDM example. These innovations and efficiency gains are crucial for practical likelihood-based inference in modern experimental settings.
@@
The remainder of this article is structured as follows. Section 2 begins with a mathematical description of the GDDM. Section 3 summarizes previous work related to this paper, including other efforts to address the computational challenges in first passage time models. In Section 4, we introduce our proposed method for computing the FPTDs of certain GDDMs, starting with analytical results for simple DDMs, then combining them sequentially based on Markovian properties. We also apply variable transformations and piecewise linear approximations to extend the method to more complex GDDMs, including those that allow for position-dependent drift and diffusion terms as well as dynamic decision boundaries [usher2001time, fengler2022, GingJehli2025]. Section 5 shows several numerical examples that demonstrate the speed and accuracy of our method for both likelihood computation and statistical inference tasks. Section 6 summarizes the present work, including its applicability and suggestions for future work. It is worth mentioning that while our examples are drawn from cognitive neuroscience and mathematical psychology, the method we propose is equally applicable to first passage time models in other fields. For instance, it could be extended to financial applications such as pricing double barrier options [lin1998double], as well as other areas where first passage times play a critical role.
@@ 2 Drift Diffusion Model
as the first passage time densities (FPTDs) on the upper boundary and lower boundary [^2] Note that “first passage time density on the upper (lower) boundary” is sometimes interpreted as the density of $\tau_{u}=\inf\{t>0:X(t)\geq u(t)\}(\tau_{\ell}=\inf\{t>0:X(t)\leq\ell(t)\})$ when only considering one upper boundary $u(t)$ (lower boundary $\ell(t)$). In this case $\tau=\tau_{u}\wedge\tau_{\ell}$. These quantities represent different aspects compared to our definition of the FPTD on the upper and lower boundaries and should not be confused., respectively, implicitly assuming throughout sufficient regularity conditions on the model for these densities to exist (see the sufficient conditions for density existence in [ferebee1982tangent, strassen1967almost]). The marginal density of $\tau$ is thus given by $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\mathbb{P}(\tau\leq t)=f_{u}(t)+f_{\ell}(t)$. With this notation, we can write the joint probability density of $(\tau,C)$ as
@@
As mentioned in the Introduction, we are motivated by practical problems for which we need to evaluate many different joint densities (3). We define $\boldsymbol{\Xi}=\{\mu(\cdot),\sigma(\cdot),u(\cdot),\ell(\cdot),q_{0}(\cdot)\}$ to be the configuration of a GDDM, where $q_{0}$ is a representation for the distribution of $X(0)$, usually a probability density function or a point mass at a deterministic starting point $x_{0}$. In statistical applications, $\boldsymbol{\Xi}$ will depend on unknown parameters and known covariates, some of which might vary within and across each trial. If the $i$-th trial results in the observation $(\tau_{i},C_{i})$, then the contribution to the likelihood for that trial is $f(\tau_{i},C_{i}\mid\boldsymbol{\Xi}_{i})$. This is simply the FPTD on the boundary $C_{i}$ evaluated at time $\tau_{i}$ using the GDDM configuration $\boldsymbol{\Xi}_{i}$ for the $i$-th trial. In experiments with a time limit, some trials might result in non-response, in which case the contribution to the likelihood is the NPP for that GDDM configuration, say $Q_{i}=Q(\boldsymbol{\Xi}_{i})$. Consequently, denoting $R_{i}=\mathbbm{1}\{\text{a response on trial $i$}\}$, the likelihood function is then given by [^3] The statistical model for $\tau$ here is a mixed distribution consisting of a continuous part supported on $\tau\in(0,T]$ (response) and a point mass at $\tau>T$ (no response). Mixed likelihoods of this form are common in survival analysis, for example see the Section 2.2 of [lawless2011statistical].
@@ 3 Related Work
Analytic or semi-analytic methods. For certain simple DDMs, analytic formulas for the first passage time densities or distributions can be derived, often in the form of infinite series. These methods are fast and accurate, but are typically applicable only for very specific cases. For example, they are available for the standard DDM where all parameters $\mu,\sigma,x_{0},u,\ell$ are constant [navarro2009fast, blurton2012fast, gondan2014even]. More recently, [srivastava2017martingale] used martingale theory to derive closed-form formulas that extend certain theoretical analyses to GDDMs.
@@
The approach in this paper is based on numerically extending these fast special cases to much broader classes of GDDMs. Along this line, earlier works [wang1997boundary, novikov1999approximations, potzelberger2001boundary, wang2007crossing] computed NPPs, while [jin2017first], [ichiba2011efficient], and the present work target FPTDs, although [jin2017first] and [ichiba2011efficient] restrict attention to the special case of a one-sided boundary. The FPTD case is more challenging but also more useful for statistical inference since it directly gives the likelihood. A key limitation of many prior approaches [wang1997boundary, potzelberger2001boundary, wang2007crossing, ichiba2011efficient, jin2017first], is that they rely on Monte Carlo integration, which is infeasible for large-scale real-world applications.
@@
Numerical solutions of governing equations. These methods exploit the known connections between diffusion processes and partial differential equations (PDEs), called Kolmogorov equations (see the Chapter 8 of [weinan2021applied] for an introduction, and Chapter 5.7 of [karatzas2014brownian] for a mathematically rigorous treatment). Both the Kolmogorov forward equation (KFE, also known as Fokker-Planck equation in physics) [shinn2020flexible, diederich2003simple] and the Kolmogorov backward equation (KBE) [voss2008fast, voss2007fast] have been exploited to derive the FPTDs. Moreover, the first passage time problem can be formulated into Volterra integral equations of renewal type [smith2000stochastic, peskir2002integral], where applications to DDMs [smith2022modeling, paninski2008integral] have also been explored. In these works, numerical methods, particularly finite difference methods [thomas2013numerical], have been developed to solve these equations, enabling the efficient computation of FPTDs in certain cases. Other numerical methods like discontinuous Galerkin methods have also been considered, see [patie2008first] and references therein. However, these approaches remain computationally expensive, especially in settings where the GDDM configuration differs across trials, including but not limited to cases with trial-specific covariates that vary dynamically over time. For a comprehensive review of these methods, we refer the readers to [richter2023diffusion].
@@
Stochastic simulation. Methods that can sample from the hitting time distribution of the target GDDM can be used to approximate FPTDs via density estimation. Efficient methods for exact sampling are challenging; see [herrmann2020exact, herrmann2022exact] and the references therein. If the goal is statistical inference, then these sampling algorithms can be incorporated into a variety of simulation-based inference (SBI) procedures; see [cranmer2020frontier] for review and [fengler2021likelihood] for an SBI example with DDMs. Nevertheless, the Monte Carlo methods underlying SBI are computationally intensive and can become impractical for large datasets, particularly when $\Xi$ involves high-dimensional parameters or covariates that vary dynamically over time and across trials.
@@ 4.1 Single-stage Model
[hall1997distribution] derives the following analytical formulas [^4] Another work along this direction is [anderson1960modification]. We adopt the results in [hall1997distribution] since they are simpler and were shown to agree with those of [anderson1960modification]. Equation (4.1), (4.1), and (4.1) have been slightly rewritten for use in this paper, compared to their original form in [hall1997distribution]. for FPTDs and NPDs in the single-stage model for the special case when $x_{0}=0$ and $\sigma=1$: We define $\overline{a}=\frac{a_{1}+a_{2}}{2},\ \overline{b}=\frac{b_{1}+b_{2}}{2},\ c=a_{1}-a_{2},\ b=\frac{b_{2}-b_{1}}{2}$, then
@@ 4.5 Transformations
A particularly simple $\widetilde{X}$ is standard Brownian motion, perhaps with a random starting point, i.e., equation (2) with $\mu(x,t)\equiv 0$ and $\sigma(x,t)\equiv 1$. A wide range of commonly used SDEs can be transformed into standard Brownian motion in this way, including geometric Brownian motion, the Ornstein-Uhlenbeck (O-U) process, and the Brownian bridge, among others. The question of whether a diffusion process can be transformed to a Brownian motion through transformation of the type in equation (13) was raised by Kolmogorov in [kolmogoroff1931analytischen]. Cherkasov gave a constructive proof of the necessary and sufficient condition for the existence of such a transformation in [cherkasov1957transformation], which is now commonly referred as the Cherkasov condition in the literature. This was recast into a simpler form in [ricciardi1976transformation], and was later used in [ricciardi1983note] to obtain the one-sided first passage time density through the integral equation method. We present the Cherkasov condition in [ricciardi1976transformation] in Appendix A.2, and instead of focusing on the condition itself, we provide specific examples of the transformation from diffusion processes to Brownian motion that are common in practical applications.
@@
where $\theta\neq 0,\lambda\in\mathbb{R},\sigma>0$ are constants. When $\theta>0$ it models the “mean-reverting” phenomenon as the process tends towards $\lambda$ over time, and it will converge to its stationary distribution $\mathcal{N}(\lambda,\frac{\sigma^{2}}{2\theta})$ as $t$ goes to infinity. This process is used, for instance, in physics to model the velocity of a particle under friction, in finance to capture mean-reverting behavior of interest rates and volatility [gardiner2009stochastic], and in cognitive science to model evidence accumulation and neural dynamics [usher2001time, Brunton2013].
The solution of (18) is given by [gardiner2009stochastic]
@@
The stochastic integral can be viewed as a time-changed Brownian motion [doob1942brownian]: